# Charged excitations made neutral: N-centered ensemble density functional theory of Fukui functions

### v1

@@ Abstract
Figure 5: Fukui functions in strong correlation regime for non-zero weights (parameters in the red inset). We compare the Padé EDFA (green) to EEXX prediction (blue). Exact results (orange) and neglect of the weight-derivative terms (grey) are plotted for reference.
This EDFA performs poorly (not shown) as its discontinuous behavior (through the sup function) and density-derivative discontinuity in |n1||n-1| means it lacks cross-derivatives (in $n$ and $\xi_{\pm}$) contributions to $\frac{\partial v_{\rm Hxc}^{\bm{\xi}}[n]}{\partial\xi_{\pm}}\Big|_{n=n^{\bm{\xi}}}$ that are important even in the high correlation regime. Even if Eq. (23) is formally exact in the strictly-correlated regime ($U\rightarrow\infty$), it is not representative of the strongly-correlated regime (typically $U\in[5,10]$). To remedy this, the sharp sup function and absolute value were each replaced by a smoothing function whose stiffness parameters increase with $U$, thus yielding cross derivative terms while still being compatible with the exact $U/t\rightarrow\infty$ limit. To set the stiffness parameters in a way consistent with what could be done for a real system, we only considered the symmetric case of 2-electrons Hubbard dimer, i.e. $\Delta v=0$ and $\bm{\xi}=0$, which is the analog of the uniform electron gas in this context. The stiffness parameters were set so that our Padé approximation reproduces best the (ground-state) Hxc kernel. The obtained EDFA is in principle applicable for any choice of weights, even if it was smoothed by learning from the ground-state symmetric case. It is thus most interesting to test its performance for non-zero weights.
[…]
This trend carries over to higher correlation, as shown on Fig. 6 for $U/t=10$ and $\xi=0.2$, which is remarkable as the stifness parameters smoothing the $U\rightarrow\infty$ discontinuities in $\gamma^{\bm{\xi}}[n]$ where fixed using only the ground-state, symmetric limit of the Hxc functional.
Figure 6: Fukui functions in very strong correlation regime for non-zero ensemble weights (parameters in the red inset). See Fig. 5 for labels.
Summary and outlook– After introducing the (in principle) exact expression of Fukui functions within Nc EDFT, we discussed the equivalence between weight-derivatives of the Hxc potential and derivative discontinuities in the Hxc kernel. The exact relations introduced by the Authors between individual true physical and non-interacting KS densitiesFromager (2025) and responsesDupuy and Fromager (2025) are thus a promising framework to access properties beyond excitation energies within EDFT.
By considering the conventional $N$-electron ground-state DFT limit of the theory, we showed how dressing the regular (ground-state) Hxc functional with a weight-dependent scaling function allows to recycle practical (semi-)local DFAs while accurately accounting for the ensemble Hxc potential weight-derivatives term, as illustrated on Fukui functions of the Hubbard Dimer. It was instrumental to the performance of this strategy that separate scaling functions would be applied to the Hx and correlation parts of the ground-state potential, as their ensemble analogs had starkly different dependence on ensemble weights and density. This is in line with analysis of low- and high-correlation limits of EDFTGould et al. (2023), predicting a dominant dependence on the ensemble density rather than on the weights in the latter case. When used properly, the scaling function approach extended the domain of applicability of its parent approximation by exploiting an accurate ground-state functional.
We also showed building a full ensemble DFA by interpolating between known low- and high-correlation limits performed very well, provided one paid attention to avoid discontinuous behavior in the functional. Smoothing over these discontinuities by learning from the symmetric ground state proved very effective, echoing recent work on the uniform electron gas as a basis to develop excited state density functionalsLoos (2025). These findings will foster DFA development in realistic systems. As additional strategies, one could also include exact constraints satisfied by ensembles in the design of scaling functions, such as the invariance of individual-state physical densities with respect to weights (see Eq. (8)), which is currently explored. As our formalism applies to any (neutral or charged) excitation, it can bring insights into processes that PPLB could not grasp, notably excited state reactivity through the relevant Fukui functionsAn et al. (2024). This is left for future work.
Acknowledgements - This work has benefited from support provided by the University of Strasbourg Institute for Advanced Study (USIAS) for a Fellowship, within the French national programme “Investment for the future” (IdEx-Unistra).

### v2

@@ Abstract
Figure S1: Fukui functions in weak correlation regime for null ensemble weights (parameters in the red inset). Full and dashed lines are $f_{+}$ and $f_{-}$ calculations respectively. Exact results (yellow lines) are compared to EEXX (blue lines) and to the neglect of weight-derivative contributions (grey lines).
The situation changes when increasing correlation up to $U=1.5$ (keeping 0 weights) as shown on Fig. S2. A noticeable discrepancy between EEXX and exact results appears for low external potentials $\Delta v<2U$, where the correlation energy is the highest. Note however the improvement over neglect of weight-derivative terms is still noticeable for $f_{-}$. Inclusion of correlation in the PT2 DFA allows to recover a close agreement with exact results in all the $\Delta v$ range.
Figure S2: Fukui functions in the low correlation regime for null ensemble weights (parameters in the red inset). We compare EEXX ensemble DFA (blue) to PT2 (green) and neglect of weight derivative terms (grey). Exact results are shown in orange.
S-II.3 Scaling functions
Inadequacy of an Hxc energy-based scaling function – Using the ratio of Hxc ensemble and ground-state energies at the ensemble density is a poor strategy to predict the Fukui function.
$s(\bm{\xi})=\frac{E^{\bm{\xi}}_{\rm Hxc}[n^{\bm{\xi}}]}{E_{\rm Hxc}[n^{\bm{\xi}}]}$ (S77)
The reason is that they encode widely-different weight-dependence, as seen on Fig. S3 where we compare the two definitions for the scaling function at $U/t=2.5$. The weight $\xi_{-}$ is kept fixed at 0, while $\xi_{+}$ is varied. Different values of the external potential, producing different ensemble densities, are considered. In all cases, it is seen that the scaling functions using energies and potentials vary in an almost opposite way, making the former unsuited to encode weight-dependence of the Hxc potential.
Figure S3: Profiles of the scaling function as a function of $\xi_{+}$ defined from the Hxc energy ratio (full lines) compared to the ratio of potentials (dashed lines). The weight $\xi_{-}$ is kept at 0. Three values of the external potential are considered.
Improvement brought by double scaling at the PT2 level – To understand the impact of the double-scaling, we compare on Fig. S4 the ensemble correlation potential predicted by the full-PT2 DFA to the one obtained by scaling the ground-state correlation potential using to the ratio of PT2 ensemble and ground-state correlation potentials. Each of the four panels correspond to a different correlation strength $U/t$. On each panel, in the 0-weight case, the scaling strategy is exact by construction, while the full PT2 correlation potential exhibits increasing errors with $U/t$ when the system is strongly asymmetric. Results for $\xi=\xi_{+}=\xi_{-}=0.05$ show the scaled potential matches closely the exact ensemble one, even for large $U/t$, with the brunt of the error around the symmetric case. This shows the weight-derivatives of the Hxc potential will be modeled more accurately for large $\Delta v$ thanks to the scaling approach, while they will not improve nor deteriorate results around $\Delta v=0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure S4: Ensemble correlation energy as a function of ensemble density. Exact results (full lines) are compared to full PT2 potential (long-dashed lines) and to the PT2 double scaling function applied to the ground state potential (short-dashed lines). Two values are considered for the weights, taken to be equal $\xi=\xi_{+}=\xi_{-}$. Each panel corresponds to a specific value of $U$.
This is illustrated on Fig. S5 for $U/t=1.5$, comparing exact weight-derivatives of the ensemble Hxc energy to the full PT2 EDFA and the PT2-derived double scaling.
Figure S5: Weight-derivatives of the ensemble Hxc potential in the 0-weight limit. Exact results (full lines) are compared to full PT2 EDFA (long-dashed lines) and PT2 scaling of ground state DFA (short-dashed lines).
S-II.4 Smoothing of Pade interpolant
Although $U\,\gamma^{\bm{\xi}}(n)$ defined through Eq. (S71) is the exact limit of $E^{\bm{\xi}}_{\rm Hxc}(n)$ for $u\rightarrow\infty$, its discontinuous behavior is a poor approximation of the strong correlation regime (say, $U/t>5$.) On Figs. S6 to S9, the right panels show it does not reproduce Fukui functions accurately for $\Delta v<U$, considering to sets of weights and two values of $U$. The discontinuities in $\gamma^{\bm{\xi}}$ derivatives are reflected in the Fukui functions, at $\Delta v=0$ for the 0-weight case, and displaced around $2\xi\,U$ for $\xi=\xi_{+}=\xi_{-}=0.2$. The left panels show the Padé approximant error comes from a missing contribution to the weight-derivative of the Hxc potential for low $\Delta v$. Its prediction for the Hxc kernel, shown in middle panels, is also poor in this range. It is reasonable to assume that for high-but-finite $U$, the Hxc energy has continuous derivatives with respect to the density and weights, which yield cross-derivative contributions to $\frac{\partial v^{\bm{\xi}}_{\rm Hxc}(n)}{\partial\xi_{\pm}}|_{n=n^{\bm{\xi}}}$. Similarly, it would induce important variations of the kernel around the derivative discontinuities of the Padé approximant that we observe.
We thus replace the original expression for $\gamma^{\bm{\xi}}(n)$ by a smoothed version:
$\begin{split}\bar{\gamma}^{\bm{\xi}}(n)&=\eta(n-1)+\frac{1}{2}\left(\xi_{+}-\xi_{-}+\frac{1}{k_{\xi}}\ell n(1+e^{k_{\xi}[\xi_{+}+\xi_{-}-2*\eta(n-1)]}\right)\\ \eta(n-1)&=\frac{2}{k_{n}}\ell n\left(1+e^{k_{n}*(n-1)}\right)-(n-1)\end{split}$ (S78)
where the parameters $k_{n}$ and $k_{\xi}$ control the stiffness of the function, and should be functions of $U$. This is an opportunity to explore how conditions on the functional in familiar regimes can guide DFA development. To do so, we set $k_{n}$ and $k_{\xi}$ so that $f^{\bm{\xi}}_{\rm Hxc}(n)$ is correctly reproduced in the symmetric case ($\Delta v=0$), corresponding to the uniform electron gas limit, for zero weights. This is illustrated on middle panels of Figs. S6 and S8. Multiple couple values $(k_{n},k_{\xi})$ can satisfy this requirement. As an additional constraint, we minimize the difference between the exact and Padé slopes of $\frac{\partial\Delta v^{\bm{\xi}}_{\rm Hxc}(n)}{\partial\xi_{\pm}}$ for $\Delta v=0$ and zero weights, as seen on left panels of aforementioned figures. This yields the parameters $(k_{n}=64,k_{\xi}=15)$ for $U=5$ and $(k_{n}=130,k_{\xi}=25)$ for $U=10$. The improvement over the Padé approximation with discontinuous derivatives can be seen on right panels of Figs. S6 and S8 for the prediction of Fukui functions. More interesting, in our opinion, is the impact outside the 0-weight limit. One can see the smoothed Padé approximant improves substantially the Hxc kernel, weight-derivatives of Hxc potential, and resulting Fukui functions for $\xi=\xi_{+}=\xi_{-}=0.2$ on Figs.S7 and S9. This is a promising result for transferability as only the 0-weight limit was used to improve the Padé approximant. Of course, other constraints could be considered, and our choice of smoothing functions is entirely arbitrary. In that regard, performing a Taylor expansion in $1/u$ around $1/u=0$, in the spirit of ISIGould et al. (2023) would be informative in realistic systems. The present constructions simply serves to demonstrate care must be taken with discontinuous derivatives in ensemble DFAs.
Refer to caption Refer to caption Refer to caption
Figure S6: Effect of smoothing the strictly-correlated limit in the Padé approximant. Left panel: Weight-derivatives of the Hxc potential. Middle panel: Hxc kernel. Right panel: Fukui functions.
Refer to caption Refer to caption Refer to caption
Figure S7: Same as Fig. S6 for non-zero weights $\xi=\xi_{+}=\xi_{-}$.
Refer to caption Refer to caption Refer to caption
Figure S8: Same as Fig. S6 with $U/t=10$.
Refer to caption Refer to caption Refer to caption
Figure S9: Same as Fig. S6 with $U/t=10$ and $\xi=\xi_{+}=\xi_{-}=0.2$.